# When RL Meets Adaptive Speculative Training:

### v1

@@
Abstract Speculative decoding can significantly accelerate LLM serving, yet most deployments today disentangle speculator training from serving, treating speculator training as a standalone offline modeling problem. We show that this decoupled formulation introduces substantial deployment and adaptation lag: (1) high time-to-serve, since a speculator must be trained offline for a considerable period before deployment; (2) delayed utility feedback, since the true end-to-end decoding speedup is only known after training and cannot be inferred reliably from acceptance rate alone due to model-architecture and system-level overheads; and (3) domain-drift degradation, as the target model is repurposed to new domains and the speculator becomes stale and less effective. To address these issues, we present Aurora, a unified training–serving system that closes the loop by continuously learning a speculator directly from live inference traces. Aurora reframes online speculator learning as an asynchronous reinforcement-learning problem: accepted tokens provide positive feedback, while rejected speculator proposals provide implicit negative feedback that we exploit to improve sample efficiency. Our design integrates an SGLang-based inference server with an asynchronous training server, enabling hot-swapped speculator updates without service interruption. Crucially, Aurora supports day-0 deployment: a speculator can be served immediately and rapidly adapted to live traffic, improving system performance while providing immediate utility feedback. Across experiments, Aurora achieves a 1.5× day-0 speedup on recently released frontier models (e.g., MiniMax M2.1 229B and Qwen3-Coder-Next 80B). Aurora also adapts effectively to distribution shifts in user traffic, delivering an additional 1.25× speedup over a well-trained but static speculator on widely used models (e.g., Qwen3 and Llama3). 00footnotetext: The project leads are Xiaoxia Wu, Chengfeng Xu and Junxiong Wang  Website: https://aurora-spec-ai.github.io/
@@ 1 Introduction
The deployment of large language models (LLM) is increasingly strained by inference costs [7]. Speculative decoding [13, 14, 16, 4, 11] is a compelling system lever: a lightweight drafter proposes multiple tokens and a high-quality verifier (the target model) checks and accepts a prefix. This reduces the number of expensive target-model decode steps while preserving output quality [18, 1, 5]. When the drafter’s proposals align closely with the verifier, speculative decoding can deliver substantial speedups at the production scale [29, 33].
@@
Most speculative decoding deployments today follow a two-stage pipeline (i) an offline training workflow that trains a draft model to maximize acceptance via supervised targets and/or off-policy distillation from target-model activations [14, 16] and (ii) a separate serving stack that loads the trained drafter and runs speculative decoding online. While this separation is organizationally convenient, we argue that it is also a key reason speculative decoding can under-deliver in real deployments. Specifically, the two-system pipeline introduces systematic challenges:
(1) Day-0 support is difficult. In practice, operators want inference speedups from day 0 of a new model release. However, two-stage speculative decoding systems typically require a lengthy calibration and training pipeline for the drafter before they deliver reliable gains. This workflow adds both wall-clock delay and engineering overhead, resulting in a high time-to-serve.
@@
(2) Training–serving mismatch causes delayed utility feedback. Offline drafter training can maximize acceptance in a controlled setting, but production speedups are ultimately determined by the deployment stack such as kernel implementations, numeric precision (FP8/FP4), batching and scheduling behavior, and even micro-architectural effects. In practice, the drafter and verifier often run under different runtime constraints, and the drafter may even come from a different model family. Such mismatches can introduce incompatibilities, non-trivial numerical drift, and unexpected throughput regressions. As a result, the true end-to-end speedup is hard to predict without serving-time evaluation.
(3) Moving-target gap from verifier drift. In real deployments, the target is not static: verifier models are frequently updated for quality, safety, cost, or hardware migration. However, the drafter is typically refreshed on a slower cadence due to retraining cost and pipeline complexity. This creates staleness: the deployed drafter tracks an outdated verifier distribution—leading to degraded speculative performance over time.
@@
(4) High infrastructure cost. Drafter-model calibration pipelines often require collecting large volumes of target model’s activations or other signals by running the target model forward on a distillation corpus. These intermediate artifacts must be stored, versioned, transferred (often across regions or between serving and training clusters), and replayed for training. At the production scale, the storage footprint can reach petabyte-level magnitude, with high cost in memory, bandwidth, and operational complexity.
@@
These gaps suggest that speculative decoding should not be treated merely as a modeling problem (“train a better drafter”), but as a joint learning-and-serving problem. Inspired by modern reinforcement learning (RL) systems that tightly integrate online inference with continuous training, we propose an asynchronous unified training-service system with speculative decoding, called Aurora (Figure 1) that co-designs drafter learning and inference within a single integrated system.

### v2

@@
Abstract Speculative decoding can significantly accelerate LLM serving, yet most deployments today disentangle speculator training from serving, treating speculator training as a standalone offline modeling problem. We show that this decoupled formulation introduces substantial deployment and adaptation lag: (1) high time-to-serve, since a speculator must be trained offline for a considerable period before deployment; (2) delayed utility feedback, since the true end-to-end decoding speedup is only known after training and cannot be inferred reliably from acceptance rate alone due to model-architecture, diverse prompt engineering, and system-level overheads; and (3) domain-drift degradation, as the target model is repurposed to new domains and the speculator becomes stale and less effective. To address these issues, we present Aurora, a unified training–serving system that closes the loop by continuously learning a speculator directly from live inference traces. Aurora reframes online speculator learning as an asynchronous reinforcement-learning problem: accepted tokens provide positive feedback, while rejected speculator proposals provide implicit negative feedback that inherits the online traffic failure signals. Our design integrates an SGLang-based inference server with an asynchronous training server, enabling hot-swapped speculator updates without service interruption. Crucially, Aurora supports day-0 deployment: a speculator can be served immediately and rapidly adapted to live traffic, improving system performance while providing immediate utility feedback. Across experiments, Aurora achieves a 1.5× day-0 speedup on recently released frontier models (e.g., MiniMax M2.1 229B and Qwen3-Coder-Next 80B). Aurora also adapts effectively to distribution shifts in user traffic, delivering an additional 1.25× speedup over a well-trained but static speculator on widely used models (e.g., Qwen3 and Llama3). 00footnotetext: The project leads are Xiaoxia Wu, Chengfeng Xu and Junxiong Wang  Website: https://aurora-spec-ai.github.io/
@@ 1 Introduction
The deployment of large language models (LLM) is increasingly strained by inference costs [2, 22, 1]. Speculative decoding [14, 15, 17, 6, 12] is a compelling system lever: a lightweight drafter proposes multiple tokens and a high-quality verifier (the target model) checks and accepts a prefix. This reduces the number of expensive target-model decode steps while preserving output quality [19, 3, 7]. When the drafter’s proposals align closely with the verifier, speculative decoding can deliver substantial speedups at the production scale [32, 36].
@@
Most speculative decoding deployments today follow a two-stage pipeline: (i) an offline training workflow that trains a drafter to maximize token acceptance via supervised targets and/or off-policy distillation from target-model activations [15, 17], and (ii) a separate serving stack that loads the trained drafter and performs speculative decoding online. While this separation is organizationally convenient, we argue that under real-world deployment constraints, speculative decoding must be reconsidered jointly from both algorithmic and system perspectives.
1.1 Algorithm gap: Speculative training should pivot toward modeling local traffic
Prior work implicitly formulates speculative training as large-scale model distillation: the drafter is optimized to approximate the target model over a large, static, and global offline data distribution. In this view, speculative decoding becomes a miniature replica of large-model pretraining: the target model fits massive datasets, while the drafter fits the target’s behavioral traces (e.g., logits or activations) across the same broad distribution.
However, this algorithm abstraction mismatches real-world serving. In deployment, inference traffic is neither global nor stationary, (1) production systems operate under local time-varying traffic distributions, and performance is determined by the current request stream, not the full behavioral manifold of the target model. A drafter that is globally optimal with respect to the target model’s distribution may be suboptimal or even inefficient under the specific workload encountered at serving time. (2) Even worse, in practical deployments, the target model itself is not static, it may frequently be updated for quality, safety, cost optimization, or hardware migration. In contrast, the drafter is typically refreshed on a much slower cadence due to retraining cost and pipeline complexity. This creates distributional staleness: the deployed drafter tracks an outdated verifier distribution, leading to degraded acceptance rates and diminishing speculative gains over time. Motivated by this gap, we propose to fundamentally redefine speculative training for real-world settings:
Instead of pursuing global optimality over the target model’s full behavioral space, we shift the objective toward rapid adaptation to a locally optimal drafter under the current traffic distribution.
Under this formulation, speculative training becomes a dynamic systems problem, in which learning objectives are tightly coupled with serving-time workload characteristics and deployment constraints.
1.2 System gap: Speculative training should be treated as a dynamic system
Despite this need for adaptivity, existing speculative training pipelines remain fundamentally static. This mismatch also introduces several system-level frictions in real-world deployments:
(1) Day-0 support is difficult.
New frontier models are released on a weekly cadence [27], and operators expect inference acceleration from day 0 of deployment. However, static speculative systems typically require lengthy calibration and retraining cycles for the drafter before delivering reliable gains. This introduces substantial wall-clock delay and engineering overhead, increasing time-to-serve and limiting operational agility.
(2) Training–serving mismatch delays utility feedback.
@@ 1.2 System gap: Speculative training should be treated as a dynamic system
Offline drafter training may maximize token acceptance in isolation, yet production speedups are ultimately determined by the entire deployment stack, including kernel implementations, numeric precision (e.g., FP8/FP4), batching and scheduling policies, prompt diversity, and even micro-architectural effects. In practice, the drafter and verifier frequently operate under different runtime constraints and may even originate from different model families. Such mismatches introduce incompatibilities, distribution drift, and unexpected throughput regressions. As a result, true end-to-end speedup cannot be reliably predicted without serving-time evaluation.
(3) High infrastructure cost.
@@
Drafter calibration pipelines often require collecting large volumes of target-model activations or related signals by executing the verifier over a distillation corpus. These intermediate artifacts must be stored, versioned, transferred (often across regions or clusters), and replayed during training. At production scale, the storage footprint can reach petabyte-level magnitude, incurring significant memory, bandwidth, and operational complexity costs.
1.3 Aurora: A unified training-serving speculative system
@@
These joint algorithm-system gaps suggest that speculative decoding should not be treated merely as a modeling problem (“train a better drafter”), but as a joint learning-and-serving problem. Inspired by modern reinforcement learning (RL) systems that tightly integrate online inference with continuous training, we propose an asynchronous unified training-service system with speculative decoding, called Aurora (Figure 1) that co-designs drafter learning and inference within a single integrated system.